In a solid, it is possible to generate sound waves with different velocities dependent on the deformation mode Sound generating volumetric deformations (compression) and shear deformations (shearing) are caled pre Songtudu waves) and shear waves (transverse waves), respectively. The measured speed of longitudinal sound waves in one type of steel is about 5930 meters per second. What is this speed in kilometers per hour (km/h?)


1
Expert's answer
2022-04-28T13:01:37-0400

"5930~\frac ms=5930\cdot \frac{3600}{1000}~\frac{km}h=21348~\frac{km}h."